3.1 Write a number sentence that will represent the structure of the following problem situations. 3.1.1 Nare had some marbles. After receiving 24 from his sister, he now has 52 marbles altogether. How many marbles did Nare have at the beginning? (2) 3.1.2 Gabriel and Miranda have R257 between them. Gabriel has R165. How much does Miranda have? (3.1.3 Shaheed bought 3 chickens for R94 each from a local supermarket. She also bought a bag of potatoes. Shaheed paid R282 in total. How much did the potatoes cost? (2) 3.1.4 Phumzile is 6 years younger than her brother, Piet. How old is Piet when Phumzile is 25 years old? (2) 3.1.5 The young ostrich, Lenka, has to take 3 steps to keep up with every step taken by her mother. How many steps must Lenka take when her mother takes 6 steps? (2) 3.1.6 Mashudu had R500 in his wallet. After buying some dinner for him and a friend, he now has R220. How much did Mashudu pay for the dinner? (2) 3.2 Name the problem type for each of the problem situations in 3.1.

1
Expert's answer
2022-05-24T17:39:44-0400

3.1.1

Nare had an unknown amount (X) of marble stones. he received 24 marbles from his sister (X + 24). After that, the number of marbles was 52 (X + 24 = 52). first we find how many marbles there were. for this we substrac the marbles we received from his sister from the total marbles:

X= 52 - 24 = 28

ANSWER: 28

Problem type: Find an unknown additive

3.1.2.

Miranda has an unknown amount of money (X). the Gabriel has R165. both have a total R257 (X + 165 = 257). find out how much money is in the mirand. for this we subtract the money in the Gabriel from the total money:

X= 257 - 165 = 92

ANSWER: Miranda had R92.

Problem type: Find an unknown additive

3.1.3.

Shaheed bought 3 chickens (94 * 3 = 282) each, R94, and a bag of potatoes (X). Spent a total of R282 for 3 chickens (R282) and potatoes. find out how much money he spent on potatoes. for this we subtract the money spent for the chicken from the total money spent:

X= 282 - 282 = 0

ANSWER: he did not spend money on potatoes.

Problem type: Find an unknown additive

3.1.4

Phumzile is 6 years younger than her brother Piet and Phumzile is 25 years old. To find out how old Piet is, we add 6 to Phumzile's age:

X= 6 + 25 = 31

ANSWER: Piet is 31 years old.

Problem type: find the sum

3.1.5

Lenka has to take 3 steps when her mother takes 1 step to reach her mother. that is, it takes 3 times as many steps. To find out how many steps Lenka takes when her mother takes 6 steps, we multiply her mother's steps by 3:

X= 6 * 3 = 18

ANSWER: 18 step.

Problem type: find the multiplier

3.1.6

Mashudu ​​had R500. he ate with a friend. for both he paid (X). after that he had R220 left. find out how much money he spent on food. for this we subtract the last money from the first money:

X= 500 - 220 = 280

ANSWER: R280 pay for the dinner.

Problem type: find the difference